OpenStudy (anonymous):

urgent please if somebody can help me with this equation (1) sin(α - β) cos(α + β) = sin α cos α - sin β cos β ) thank you

OpenStudy (campbell_st):

well you need to use the expansions sin(A - B) = sin(A) cos(B) - cos(A)sin(B) and then cos(A + B) = cos(A) cos(B) - sin(A) sin(B) multiply the 2 together and you'll get the right hand side after collecting like terms.

OpenStudy (anonymous):

thanks cmapbell_st